OpenStudy (anonymous):

(5/x-2) - (3/x-2) a. 2x-16/(x+2)(x-2) b. 2/(x+2)(x-2) c. 2x-4/(x+2)(x-2) d. 2x+16/(x+2)(x-2)

OpenStudy (anonymous):

\[\frac{ 5 }{ x-2 }-\frac{ 3 }{ x-2 }\] is that the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok if you subtract them what do you get?

OpenStudy (anonymous):

are the denominators both x-2?

OpenStudy (anonymous):

\[\frac{ a }{ b }-\frac{ c }{ b }=\frac{ a-c }{ b }\]

OpenStudy (anonymous):

@alyssababy7 we will help you get the answer if you try

OpenStudy (anonymous):

none of the answers you provide fit the problem as described. However. if the problem is: (5/(x+2)) - (3/(x-2) then we can work something out

OpenStudy (anonymous):

i think B

OpenStudy (anonymous):

@gerryliyana its not; please dont give answers, we want to help her get the answer not just give it to her

OpenStudy (anonymous):

ok...,